OpenStudy (anonymous):

What is the solution set of |x – 4| + 7 = 4? What is the solution set of |2x – 3| – 10 = –1? What are the anwsers to these 2 problems & how do you get em?

hartnn (hartnn):

1. for 1st problem, subtract 7 from both sides, what u get ?

OpenStudy (anonymous):

-(x-4) +7 = 4 and (x-4) + 7 = 4 ... thats for the first on eand then solve....

OpenStudy (anonymous):

find x for both equations adn you have your two answers... then do the same thing for the second question

OpenStudy (anonymous):

so the 1st ones no solution?

OpenStudy (anonymous):

I really have no idea about any of this stuffff?

OpenStudy (anonymous):

no there are two answers... one for each equation i just made

hartnn (hartnn):

|a|=b can be written as a=b OR a=-b and the two solutions for a are b,-b Apply same principle here.

OpenStudy (anonymous):

I'll solve the first one |x – 4| + 7 = 4? Okay, you're going to start by isolating the absolute value, so: |x-4| + 7- 7= 4 - 7 and you're left with |x-4|=-3 Okay, so consider the positive an negative cases: x-4=-3 OR x-4=3 Solve for x for each so: x-4=-3 x=1 AND x-4=3 x=7 So x=-3 and x=7

hartnn (hartnn):

lol. there is NO SOLUTION for first :P

OpenStudy (anonymous):

So the first step to solving these is to get the absolute value by itself? & wait what do u mean theres no solution for the 1st one?????

hartnn (hartnn):

@melbel u know why ?

hartnn (hartnn):

YES! get the absolute value by itself.

hartnn (hartnn):

|x-4|=-3 for first right ?

OpenStudy (anonymous):

okay cool thanks! & i do that by subtracting/multiplying..... from both side to get that? & y is #one no solution?

hartnn (hartnn):

the meaning of |a| is that a *cannot* take negative values. u have |x-4|=-3 is this possible ?

OpenStudy (anonymous):

i'd no? because there both negative but 2 negatives = a positive right?

OpenStudy (anonymous):

*say

OpenStudy (anonymous):

yes that correct

hartnn (hartnn):

i would say |..*anything*..| cannot be negative so |x-4| cannot be negative.(but here its -3) hence there is no solution for first.

OpenStudy (anonymous):

Duh, hartnn! I didn't see that. I should be beaten. Yeah, no solution! I'd like to say that it's because it's 2 AM, but no, I should I noticed that right off the bat. Hartnn is right.

OpenStudy (anonymous):

thanks for that tip. so basically when you first see a problem and u see that there is a negative in the absolute values bar then u already know that its no solution?

OpenStudy (anonymous):

No, if an absolute value (after being isolated) is equal to a negative.

hartnn (hartnn):

as u rightly mentioned, first get the absolute value by itself if its negative, then no solution.

OpenStudy (anonymous):

So |x+4| = -3 has no solution because the absolute value cannot be negative.

OpenStudy (anonymous):

But |x-4| = 3 has a solution, because the 3 is positive.

hartnn (hartnn):

go for 2nd problem now, add 10 to both sides.

OpenStudy (anonymous):

oh okay! thank yall so much! yalll have no idea how muched u just helped me! i dont get math at allll its not my subject!

OpenStudy (anonymous):

You can look at it at the very basic level of absolute values, so: What is the absolute value of: |7| it's 7 What about |-7| Still 7. You can't get a negative.

OpenStudy (anonymous):

|2x-7|=-1

hartnn (hartnn):

thats not the correct way, keep |2x-3| as it is |2x-3|-10+10=-1+10 |2x-3| +0=9 |2x-3|=9 got this ?

OpenStudy (anonymous):

|2x – 3| – 10 = –1 |2x-3| = 9 So 2x-3 = 9 OR 2x-3 = -9 First case: 2x-3=9 2x=12 x=6 Second case: 2x-3=-9 2x=-6 x=-3 So x=6 OR x=-3

OpenStudy (anonymous):

correct

hartnn (hartnn):

thats right @melbel ,but isn't it better that @CIERA1338 tries that on her own.?....

hartnn (hartnn):

no problem, i hope she doesn't *look* through your solution and try herself...

OpenStudy (anonymous):

thank you all so much! but hartnn is right! It's just hard for me to work out the problems i get so confused easily in math...

OpenStudy (anonymous):

thank you all again! (: